OpenStudy (anonymous):

evaluate the function at x=3 and x=2/ write undefined if appropriate. x-2 y= ----------- x^2+2x-3

OpenStudy (anonymous):

y(3) = 3-2/9+6-3=1/12 y(2) =2-2/4+4-3 = 0/5 =0